OpenStudy (kittiwitti1):

Quantum # Practice. A. When n=3, L (caps to minimize confusion) can have values of: for 3d orbital L has value of: B. When n=4, L can have values of: For 4p orbital, L has value of: *** Check my answers please! The principal quantum # n can have the values (1-7) The angular momentum quantum # L can have integer values from (0 to n-1) The magnetic quantum # m can have integer values from (-L to L)

OpenStudy (aaronq):

\(l=n-1\geq 0\) s corresponds to \(l=0\) p corresponds to \(l=1\) d corresponds to \(l=2\) f corresponds to \(l=3\)

OpenStudy (kittiwitti1):

I put "when n=3 then L= -3,-2,-1,0,1,2,3" ?

OpenStudy (aaronq):

nope, \(m_l=\pm l\) however \(l\) can't be negative

OpenStudy (kittiwitti1):

oh right... woops

OpenStudy (kittiwitti1):

n=3 then ... 0, 1, 2 for L

OpenStudy (kittiwitti1):

and then if it's 3d then I get L=2

OpenStudy (aaronq):

so the values you wrote are incorrect. for n=3, \(l\) can be 2, 1, or 0 and say for \(l=2\), then \(m_l=-2,-1,0,+1,+2\)

OpenStudy (aaronq):

"and then if it's 3d then I get L=2" yes this is right.

OpenStudy (kittiwitti1):

okay thank you

OpenStudy (aaronq):

no problem

OpenStudy (kittiwitti1):

so basically L is everything from 0 to "n-1"?

OpenStudy (aaronq):

Yeah, those are possible values when only the principal number is specified

OpenStudy (kittiwitti1):

okay

OpenStudy (kittiwitti1):

how do I calculate whether s is \(\pm\) though?

OpenStudy (mayankdevnani):

s orbital can't be negative

OpenStudy (mayankdevnani):

s,p,d,f,s.... they are never negative

OpenStudy (kittiwitti1):

I meant s quantum # oops sorry

OpenStudy (mayankdevnani):

O.o

OpenStudy (mayankdevnani):

wym ?

OpenStudy (kittiwitti1):

spin quantum

OpenStudy (mayankdevnani):

oh

OpenStudy (mayankdevnani):

there are two values of `s quantum` +1/2 and -1/2 +1/2 for clockwise direction -1/2 for anti clockwise direction

OpenStudy (kittiwitti1):

yes but how to calculate

OpenStudy (mayankdevnani):

well we can't calculate it

OpenStudy (mayankdevnani):

it's obvious that we have +1/2 and -1/2

OpenStudy (mayankdevnani):

hey,tell me what is the full question? we'll discuss it

OpenStudy (kittiwitti1):

it is more concepts than questions right now...

OpenStudy (mayankdevnani):

s orbitals are non directional

OpenStudy (mayankdevnani):